# KG-Infused RAG: Augmenting Corpus-Based RAG with External Knowledge Graphs

### v1

@@ 5.1 Main Results
Table 1 presents the overall performance of our method across five datasets.
Performance on Complex QA Tasks.
Our method consistently outperforms all standard baselines on four multi-hop datasets, using both Qwen2.5-7B and LLaMA3.1-8B backbones. For example, with two rounds of activation, KG-Infused RAG based on the DPO-trained LLaMA3.1-8B improves the Avg metric over Vanilla RAG by 3.8% to 13.8%. These consistent gains demonstrate the effectiveness of activation-guided KG retrieval in supporting multi-step reasoning and answer generation. In contrast, Vanilla QE does not show clear improvements over Vanilla RAG, suggesting that relying solely on LLM-based query expansion, without external structured knowledge, offers limited benefits for complex reasoning tasks. Overall, these results indicate that KG-based query expansion and knowledge augmentation play a critical role in enhancing multi-hop QA. By guiding retrieval with KG facts, our method better captures relevant context and supports more accurate reasoning.
Performance on Commonsense QA Task.
On StrategyQA, which requires implicit commonsense reasoning and binary decisions, our method still brings positive improvements over baselines. This demonstrates the generality of our activation-based retrieval, which remains effective even when explicit multi-hop reasoning is less critical.
@@
We further apply DPO to enhance the LLM’s instruction-following ability in key stages of our pipeline. Although the training data comes solely from the KG-guided KA stage, we observe consistent performance gains on both multi-hop and commonsense QA tasks. In most cases, models perform better after lightweight DPO tuning. Notably, LLaMA3.1-8B shows greater improvement than Qwen2.5-7B, with Avg gains ranging from 1.1% to 9.4% across five datasets. These results suggest that even minimal preference tuning can effectively improve the model’s ability to incorporate KG facts, leading to further enhancements in our overall framework.
@@ 5.2 Plug-in Results
We apply KG-Infused RAG as a plug-in module to the existing corpus-based RAG method Self-RAG. As shown in Table 2, integrating KG-Infused RAG leads to performance improvements on four datasets, with Avg gains ranging from 5.1% to 16.9% under the 2-round activation setting. Although the Avg score on HotpotQA slightly drops by 0.8%, the method remains effective overall.
These results validate the flexibility and generalizability of our activation-based KG retrieval module, which can enhance existing RAG pipelines. We also apply KG-Infused RAGto Self-Ask, with results presented in Appendix F.1.
@@ 5.3 Ablation Study
To evaluate the contributions of two key components in KG-Infused RAG: KG-Based QE and KG-Aug Gen, we conduct an ablation study, with results shown in Table 3.
Table 2: Plug-in experimental results (%) of Self-RAG. KG-Infused RAG as a plug-in module is implemented using the DPO-trained LLaMA3.1-8B.
Method HotpotQA 2WikiMQA MuSiQue Bamboogle StrategyQA
Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg Acc
Self-RAG (selfrag-llama2-7b) 26.2 33.7 24.0 28.0 23.0 27.3 22.0 24.1 4.4 10.4 3.8 6.2 8.8 17.3 8.0 11.4 58.0
+ KG-Infused RAG (w/ 2-Round) 26.6 32.1 22.8 27.2 35.2 39.6 30.0 34.9 9.4 15.8 8.8 11.3 27.2 32.8 24.8 28.3 63.8
+ KG-Infused RAG (w/ 3-Round) 25.8 32.7 21.6 26.7 38.8 41.0 31.4 37.1 10.8 17.4 9.0 12.4 25.6 31.9 22.4 26.6 63.4
Table 3: Ablation study results (%). Vanilla RAG results are based on LLaMA3.1-8B, while all other variants are based on the DPO-trained LLaMA3.1-8B.
Method HotpotQA 2WikiMQA MuSiQue Bamboogle StrategyQA
Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg Acc
Vanilla RAG (LLaMA3.1-8B) 32.2 39.0 28.8 33.4 30.6 31.6 25.0 29.1 3.6 9.6 2.0 5.1 20.0 24.8 18.4 21.1 66.2
+ KG-Based QE (w/ 2-Round) 36.2 42.9 31.0 36.7 34.6 36.9 28.8 33.4 9.6 15.8 6.2 10.5 25.6 30.2 22.4 26.1 69.0
+ KG-Aug Gen (w/ 2-Round) 36.0 41.9 29.8 35.9 42.6 45.8 35.4 41.3 10.8 16.2 7.0 11.3 32.8 38.5 29.6 33.6 69.0
+ Both (w/ 2-Round) 36.6 43.8 31.0 37.1 45.0 46.8 37.0 42.9 12.2 18.5 8.8 13.2 31.2 33.6 24.8 29.9 69.4
Adding either component to Vanilla RAG yields consistent improvements across all datasets, confirming their individual effectiveness. Specifically, KG-Based QE improves retrieval by activating structured knowledge to enrich queries, while KG-Aug Gen enhances factual grounding by injecting relevant triples during generation. Combining both (“+ Both”) yields the strongest or competitive results in most cases, demonstrating their complementary strengths. However, in some cases, using only KG-Based QE or KG-Aug Gen yields better performance than combining both. This suggests that the effectiveness of KG activation depends on the task: some questions benefit more from KG-enhanced retrieval, while others are better supported by KG-informed answer generation.
@@ 5.4 Analysis
+ KG-Based QE (1-Round) 45.0 40.0 19.4 24.8
+ KG-Based QE (2-Round) 47.0 41.0 21.8 24.8
We evaluate the effectiveness of KG-Based QE from a retrieval perspective by comparing Vanilla RAG with and without KG-based activation. For Vanilla RAG, the top-6 passages are retrieved using only the raw query $q$, while for Vanilla RAG + KG-Based QE, 3 passages are retrieved using $q$ and 3 using the expanded query $q^{\prime}$ enriched with KG facts.
@@
Figure 2: Impact of the number of activation rounds on KG-Infused RAG. Vanilla RAG and Vanilla QE use LLaMA3.1-8B, while others use the DPO-trained version.
We investigate how the number of activation rounds affects performance. We set a maximum number of rounds to 6 and evaluate performance as rounds increase. Here, KG-Based QE and KG-Aug Gen refer to adding each component individually on top of Vanilla RAG. As shown in Figure 2, 2WikiMQA and MuSiQue see clear gains in the second activation round, but performance fluctuates slightly afterwards. HotpotQA peaks at the first activation, followed by a slight drop. These results suggest that more activation rounds do not always improve outcomes. Limiting the number of activation rounds to one or two is often both efficient and effective.
Furthermore, we observe that for 2WikiMQA and MuSiQue, KG-Aug Gen contributes more to KG-Infused RAG than KG-Based QE. In contrast, on HotpotQA, KG-Based QE alone outperforms the full method. This indicates that KG-Infused RAG does not always outperform its individual components. In some cases, using a single module yields better performance.
@@ 6 Conclusion
In this work, we identify key limitations of existing RAG systems: single-source retrieval and and the absence of a human-like spreading activation process. To address these issues, we propose KG-Infused RAG, a framework that incorporates human-curated KGs to guide query expansion and knowledge augmentation via spreading activation. Extensive experiments show that KG-Infused RAG consistently improves QA performance and can be integrated as a plug-in to enhance corpus-based RAG methods like Self-RAG.
Appendix A Illustrations of Retrieval and Activation Processes
A.1 Comparison of Retrieval Processes
Figure 3 illustrates the differences in retrieval strategies between our KG-Infused RAG framework and existing RAG methods.
Figure 3: Comparison of retrieval processes: KG-Infused RAG vs. Existing RAG Methods.
A.2 A Case of the Accumulated Subgraph from Spreading Activation
Figure 4 illustrates a case of the accumulated subgraph $\mathcal{G}_{q}^{act}$ in Wikidata5M-KG.
Figure 4: An example of the accumulated subgraph $\mathcal{G}_{q}^{act}$ from spreading activation in Wikidata5M-KG. This case, taken from the MuSiQue dataset, illustrates the resulting subgraph $\mathcal{G}_{q}^{act}$ constructed through KG-guided spreading activation. Due to space limitations, we only show part of the accumulated subgraph, and some activated entities are omitted from the figure.
Appendix B KG Preprocessing
B.1 Preprocessing Details
@@
We filter Wikidata5M by removing 5.65% of entities that lack descriptions or associated triples, resulting in a refined KG, denoted as Wikidata5M-KG. Compared to constructing task-specific KGs via LLMs, using Wikidata5M-KG ensures higher factual accuracy and significantly reduces preprocessing overhead.
B.2 Statistics of KG
The statistics of Wikidata5M-KG are shown in Table 5. During preprocessing, some entities and relations were removed for failing to meet two constraints: Entity Description Completeness and Entity Triples Completeness. Specifically, 5.65% of entities and 2.17% of relations in the original Wikidata5M were discarded. As a result of these removals, only 1.72% of the triples were lost, indicating that the overall triple loss from the original dataset is minimal.
@@ B.2 Statistics of KG
Wikidata5M 4944931 828 21354359
Wikidata5M-KG 4665331 810 20987217
Appendix C Experimental Setup Details
C.1 Datasets & Evaluation
We summarize all datasets and the evaluation details in Table 6.
@@ C.1 Datasets & Evaluation
C.2 Passage Retrieval
Table 7 summarizes the number of retrieved passages used by each method. Notably, to mitigate duplication between passages retrieved by $q$ and $q^{\prime}$, we adopt the following strategy: (1) retrieve the top $k_{p}/2$ passages using $q$ and the top $k_{p}$ passages using $q^{\prime}$; (2) select the top $k_{p}/2$ passages retrieved using using $q^{\prime}$ that do not overlap with those retrieved using $q$; (3) merge both sets to obtain $k_{p}$ unique passages for downstream processing.
Table 7: Overview of passage retrieval configurations for all methods.
Method # Passages from $q$ Passages from $q^{\prime}$ Total
Vanilla RAG 6 / 6
Vanilla QE 3 3 6
KG-Infused RAG 3 3 6
Self-Ask 6 0 6
Self-RAG 6 0 6
Self-Ask+KG-Infused RAG 3 3 6
Self-RAG+KG-Infused RAG 3 3 6
Appendix D Prompt Details
D.1 Prompts for Inference
@@
Prompt of the Passage Note Construction Stage Instructions Based on the provided document content, write a note. The note should integrate all relevant information from the original text that can help answer the specified question and form a coherent paragraph. Please ensure that the note includes all original text information useful for answering the question. Based on the provided document content, write a note. The note should integrate all relevant information from the original text that can help answer the specified question and form a coherent paragraph. Please ensure that the note includes all original text information useful for answering the question. Question to be answered: {question} Document content: {passages} Note:
@@
D.2 Prompts for DPO Data Collection
@@ D.3 Prompts of Self-ASK
Appendix E DPO Training Details
Training Data.
We construct the DPO training set using outputs from the knowledge augmentation stage of KG-Infused RAG with two rounds of spreading activation, based on the 2WikiMultiHopQA training set, sampling 1500 examples from each round to obtain over 3000 input instances. For each input $x$, we generate 6 candidate outputs by sampling the LLM with different decoding configurations, varying both the temperature and top-$p$ parameters. Specifically, the sampling settings used are:
[…]
We then prompt GPT-4o-mini to identify the best and worst outputs among the candidates for each input. After filtering low-confidence or ambiguous judgments, we obtain 2909 preference-labeled examples for LLaMA3.1-8B and 2461 for Qwen2.5-7B, which are used to train their respective DPO models.
@@ Appendix F Additional Experimental Results
F.1 Experimental Result of Self-Ask
Applying KG-Infused RAG as a Plug-in to Self-Ask.
For the Self-Ask baseline, to ensure consistent settings and reduce API costs, we replace the original web-based search engine with our local corpus retriever and substitute the GPT-3 generator with an open-source LLM. The Self-Ask baseline generates one sub-question (“follow up” question) iteratively based on few-shot demonstrations and the previous reasoning process, then retrieves passages for the sub-question and generates a corresponding answer. This continues until the model determines it can produce the final answer. To incorporate KG-Infused RAGas a plug-in module, we enhance the sub-question answering stage. Specifically, we first perform KG-Based QE on each sub-question to retrieve new passages, followed by KG-Aug Gen to enhance the retrieved passages with relevant KG triples. Answers are then generated based on the fact-enhanced notes.
Results of Self-Ask combined with KG-Infused RAG.

### v2

@@ 5.1 Main Results
Table 3: Plug-in experimental results (%) of Self-RAG and DeepNote. KG-Infused RAG as a plug-in module is implemented using the DPO-trained LLaMA3.1-8B.
Method HotpotQA 2WikiMQA MuSiQue Bamboogle
Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg
Self-RAG (selfrag-llama2-7b) 26.2 33.7 24.0 28.0 23.0 27.3 22.0 24.1 4.4 10.4 3.8 6.2 8.8 17.3 8.0 11.4
+ KG-Infused RAG (w/ 2-Round) 26.6 32.1 22.8 27.2 35.2 39.6 30.0 34.9 9.4 15.8 8.8 11.3 27.2 32.8 24.8 28.3
DeepNote (LLaMA3.1-8B-Instruct) 41.6 49.3 37.8 42.9 38.6 37.6 28.8 35.0 10.6 16.9 7.4 11.6 32.8 40.7 29.6 34.4
+ KG-Infused RAG (w/ 2-Round) 43.4 49.0 37.6 43.3 46.2 47.5 37.0 43.6 12.8 19.0 9.4 13.7 32.8 39.8 30.4 34.3
Table 1 presents the overall performance of our method across five datasets, while Table 2 presents a comparison against other KG-based methods.
Results on QA Tasks.
As shown in Table 1, our method consistently outperforms all standard baselines across different model families and sizes on both multi-hop QA and commonsense QA tasks. For example, with two rounds of activation, it achieves improvements of up to 18.7% across LLaMA3.1 models. These results demonstrate the effectiveness of activation-guided KG retrieval in supporting multi-step reasoning. By contrast, Vanilla QE shows little benefit, underscoring the importance of structured knowledge for complex tasks. Overall, the results demonstrate that KG-based query expansion and knowledge augmentation are essential for enhancing both multi-hop and commonsense QA, enabling models to capture richer context and reason more reliably.
@@
Although the DPO training data comes from 2WikiMQA, we observe gains in both in-domain and out-of-domain settings. For example, LLaMA3.1-8B-Instruct achieves a 9.4% improvement on 2WikiMQA, while also showing smaller gains on other datasets. These results suggest that even limited preference tuning can substantially enhance the model’s ability to integrate KG facts, thereby improving the overall performance of our framework.
Comparison with KG-Based Methods.
[…]
@@ 5.2 Plug-in Results
As shown in Table 3, integrating KG-Infused RAG into Self-RAG and DeepNote equips these methods with the ability to retrieve and reason over external KG facts, thereby strengthening their multi-hop reasoning. On 2WikiMQA, for example, Self-RAG combined with our method improves the Avg score by 10.8%, while DeepNote achieves an 8.6% gain. These results demonstrate the flexibility and generalizability of our activation-based KG retrieval module, which effectively enhances non-KG-based RAG pipelines. Additional experiments with Self-Ask are reported in Appendix F.1.
@@ 5.3 Ablation Study
Table 4: Ablation study results (%). Results are based on the DPO-trained LLaMA3.1-8B-Instruct.
Method HotpotQA 2WikiMQA MuSiQue Bamboogle
Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg Acc F1 EM Avg
KG-Infused RAG (w/2-Round) 36.6 43.8 31.0 37.1 45.0 46.8 37.0 42.9 12.2 18.5 8.8 13.2 31.2 33.6 24.8 29.9
     w/o KG-Based QE 36.0 41.9 29.8 35.9 42.6 45.8 35.4 41.3 10.8 16.2 7.0 11.3 32.8 38.5 29.6 33.6
     w/o KG-Aug Gen 36.2 42.9 31.0 36.7 34.6 36.9 28.8 33.4 9.6 15.8 6.2 10.5 25.6 30.2 22.4 26.1
     w/o Both 34.6 41.8 30.6 35.7 31.6 34.2 27.4 31.1 6.4 12.6 4.0 7.7 21.6 28.8 20.0 23.5
We evaluate the effectiveness of KG-guided spreading activation by ablating its two downstream modules, KG-Based QE and KG-Aug Gen. Both depend on the constructed subgraph summary $\mathcal{S}_{\mathcal{G},q}^{act}$.
As shown in Table 4, KG-Infused RAG outperforms the variant without KG-Based QE in most cases, indicating that spreading activation enhances retrieval by injecting structured knowledge into the query, thereby improving answer generation. Likewise, removing KG-Aug Gen leads to notable performance drops, suggesting that structured knowledge from spreading activation supplements retrieved passages and improves factual grounding. The improvements, however, are not uniform and may vary with the baseline retrieval performance and the alignment between the KG and the dataset.
@@ 5.4 Analysis
+ KG-Based QE (w/1-Round) 45.0 40.0 19.4 24.8
+ KG-Based QE (w/2-Round) 47.0 41.0 21.8 24.8
We evaluate the effectiveness of KG-Based QE by comparing Vanilla RAG with and without KG-based activation. For Vanilla RAG, the top-6 passages are retrieved using only the raw query $q$. For Vanilla RAG + KG-Based QE, three passages are retrieved with $q$ and three with the expanded query $q^{\prime}$ enriched with KG facts.
@@
Figure 2: Impact of the number of activation rounds on KG-Infused RAG. Vanilla RAG and Vanilla QE use Qwen2.5-7B-Instruct, while others use the DPO-trained version.
We investigate the impact of activation depth, allowing up to six rounds, where KG-Based QE and KG-Aug Gen denote adding each component individually on top of Vanilla RAG. As shown in Figure 2, 2WikiMQA, MuSiQue, and Bamboogle achieve clear gains in the second activation round, though performance fluctuates thereafter, while HotpotQA peaks at the first round and then slightly declines. These results suggest that additional activation rounds do not guarantee improvements; limiting to one or two rounds is often both efficient and effective. Moreover, we find that KG-Aug Gen contributes more than KG-Based QE on 2WikiMQA and Bomboogle, whereas on HotpotQA, KG-Based QE alone performs better. A possible explanation is that the relative effectiveness depends on the alignment between the dataset and the KG, as well as the LLM’s ability to exploit structured knowledge.
@@ 6 Conclusion
In this work, we address limitations of existing RAG methods, which either rely solely on text corpora and overlook structural knowledge, or construct ad-hoc KGs with high cost and low reliability. To overcome these issues, we propose KG-Infused RAG, a framework that leverages pre-existing large-scale KGs and applies spreading activation across retrieval and generation. Experiments show that KG-Infused RAG achieves superior efficiency and performance compared with both vanilla RAG and KG-based methods, and it can also serve as a plug-in to strengthen corpus-based RAG methods such as Self-RAG. These results underscore both the effectiveness and flexibility of our method.
Appendix A Knowledge Graph Construction and Preprocessing
A.1 KG Construction: Data Preparation
We prepare our KG $\mathcal{G}$ by extracting and preprocessing triples from Wikidata5M (Wang et al., 2021), a large-scale knowledge graph dataset derived from Wikidata and Wikipedia. Wikidata5M spans multiple domains, making it suitable for open-domain QA tasks. The KG used in KG-Infused RAG is defined as:
where $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{D}_{e}$ denote the set of entities, relations, and entity descriptions, respectively. We preprocess Wikidata5M to meet our requirements for the KG, resulting in Wikidata5M-KG with approximately $21$ million triples.
A.2 Preprocessing Details
@@
We filter Wikidata5M by removing 5.655.65% of entities that lack descriptions or associated triples, resulting in a refined KG, denoted as Wikidata5M-KG. Compared with constructing task-specific KGs using LLMs, Wikidata5M-KG offers richer coverage (more entities and relations) and greater efficiency, as it eliminates the need for LLM-based generation.
A.3 Statistics of Wikidata5M-KG
The statistics of Wikidata5M-KG are shown in Table 6. During preprocessing, some entities and relations were removed for failing to satisfy the two constraints above. Specifically, $5.65\%$ of entities and $2.17\%$ of relations in the original Wikidata5M were discarded. As a result, only $1.72\%$ of triples were removed, indicating that the overall triple loss relative to the original dataset is minimal.
@@ A.3 Statistics of Wikidata5M-KG
Wikidata5M 4,944,931 828 21,354,359
Wikidata5M-KG 4,665,331 810 20,987,217
Appendix B Experimental Setup Details
B.1 Datasets & Evaluation
Table 7 summarizes the datasets and evaluation details.
@@ B.1 Datasets & Evaluation
B.2 Model Backbones & Retrieval Settings
Table 8 summarizes the model backbones and retrieval settings used across our experiments. For LightRAG, top-$k$ denotes the number of top items to retrieve: it represents entities in the “local” mode and relations in the “global” mode. For GraphRAG (local), the default configuration retrieves $10$ entities and $10$ relations, whereas GraphRAG (global) adopts a map–reduce strategy over all relevant communities without a fixed top-$k$ limit. For the other methods, top-$k$ refers to the number of retrieved passages or chunks.
Table 8: Overview of methods, backbones, and Top-k values.
Method Model Backbone Top-$k$
Vanilla RAG Qwen2.5 / LLaMA3.1 / GPT-4o series 6
Vanilla QE Qwen2.5 / LLaMA3.1 / GPT-4o series 6
Self-Ask Qwen2.5 6
Self-RAG Qwen2.5 / LLaMA3.1 6
DeepNote Qwen2.5 / LLaMA3.1 6
LightRAG (local) GPT-4o-mini 60
LightRAG (global) GPT-4o-mini 60
GraphRAG (local) GPT-4o-mini 20
GraphRAG (global) GPT-4o-mini /
KG2RAG LLaMA3.1 6
KG-Infused RAG Qwen2.5 / LLaMA3.1 / GPT-4o series 6
Appendix C Prompt Details
C.1 Prompts for Inference
@@
Prompt of the Passage Note Construction Stage Instructions Based on the provided document content, write a note. The note should integrate all relevant information from the original text that can help answer the specified question and form a coherent paragraph. Please ensure that the note includes all original text information useful for answering the question. Question to be answered: {question} Document content: {passages} Note:
@@
C.2 Prompts for DPO Data Collection
@@ C.3 Prompts of Self-ASK
Appendix D DPO Training Details
DPO Data Construction.
We use the LLM GPT-4o-mini to construct the dataset $D_{\mathrm{KA}}$ used for DPO. Specifically, we sample examples from the 2WikiMultiHopQA training set (Ho et al., 2020) and collect intermediate outputs from the KG-guided KA stage. For each input $x$, we generate six candidate outputs using diverse decoding strategies, varying both the temperature and top-$p$ parameters. The sampling configurations are:
$\texttt{(temperature, top-}p\texttt{)}=\{(0.0,1.0),(0.3,0.95),(0.5,0.9),(0.7,0.9),(0.9,0.8),(1.0,0.7)\}.$
We prompt GPT-4o-mini to select the best and worst candidate for each input. After filtering out low-confidence or ambiguous judgments, we construct preference-labeled triples $(x,y^{-},y^{+})$ and form the DPO dataset:
The DPO training set comprises examples from the first and second rounds of spreading activation: we sample 1,500 from each round, yielding more than 3,000 input instances. The resulting dataset is used to train DPO models for LLaMA3.1-8B-Instruct (2,909 examples) and Qwen2.5-7B-Instruct (2,461 examples). Statistics of the training data are shown in Table 20.
@@ Appendix D DPO Training Details
Appendix E Implementation Details for Plug-in Baselines
Self-RAG
For Self-RAG, we follow the official setting, in which multiple passages are retrieved once and subsequently used for short-form QA tasks. We replace the original retrieval results with those obtained via KG-Based QE after KG-Guided Spreading Activation, and then apply KG-Aug Gen to each passage individually. For generation, we employ selfrag-llama2-7b [^2] https://huggingface.co/selfrag/selfrag_llama2_7b as the generator, which is additionally fine-tuned following the original Self-RAG paper.
DeepNote
For DeepNote, since both its query refinement and our KG-Based QE serve the purpose of generating new queries, we preserve its original refinement step to maintain methodological integrity. Instead, we perform KG-Guided Spreading Activation to obtain the KG subgraph summary $\mathcal{S}_{\mathcal{G},q}^{act}$, followed by KG-Guided KA to construct a fact-enhanced note $\mathcal{S}_{q}^{\text{final}}$, which is subsequently used for answer generation.
Self-Ask.
For Self-Ask, to ensure consistent settings and reduce API costs, we replace the original web-based search engine with our local corpus retriever and substitute the GPT-3 generator with the Qwen2.5-7B-Instruct. The Self-Ask baseline generates one sub-question (“follow up” question) iteratively based on few-shot demonstrations and the previous reasoning process, then retrieves passages for the sub-question and generates a corresponding answer. This continues until the model determines it can produce the final answer. To incorporate KG-Infused RAG as a plug-in module, we enhance the sub-question answering stage. Specifically, we first perform KG-Based QE on each sub-question to retrieve new passages, followed by KG-Aug Gen to enhance the retrieved passages with relevant KG triples. Answers are then generated based on the fact-enhanced notes.
@@ Appendix F Additional Experimental Results
F.1 Results with Self-Ask